# BEVTraj: Map-Free End-to-End Trajectory Prediction in Bird’s-Eye View with Deformable Attention and Sparse Goal Proposals

### v1

@@ Abstract
In autonomous driving, trajectory prediction is essential for ensuring safe and efficient navigation. To improve prediction accuracy, recent approaches often rely on pre-built high-definition (HD) maps or real-time local map construction modules to incorporate static environmental information. However, pre-built HD maps are limited to specific regions and cannot adapt to transient changes. In addition, local map construction modules, which recognize only predefined elements, may fail to capture critical scene details or introduce errors that degrade prediction performance. To overcome these limitations, we propose Bird’s-Eye View Trajectory Prediction (BEVTraj), a novel trajectory prediction framework that operates directly in the bird’s-eye view (BEV) space utilizing real-time sensor data without relying on any pre-built maps. The BEVTraj leverages deformable attention to efficiently extract relevant context from dense BEV features. Furthermore, we introduce a Sparse Goal Candidate Proposal (SGCP) module, which enables full end-to-end prediction without requiring any post-processing steps. Extensive experiments demonstrate that the BEVTraj achieves performance comparable to state-of-the-art HD map-based models while offering greater flexibility by eliminating the dependency on pre-built maps. The source code is available at https://github.com/Kongminsang/bevtraj.
Index Terms:
Trajectory prediction, autonomous vehicles, deep learning, sensor fusion, deformable attention, bird’s-eye-view.
@@ I Introduction
Trajectory prediction plays a pivotal role in autonomous driving, enabling vehicles to anticipate the future movements of surrounding agents and make safer, more informed decisions. However, this task remains inherently difficult due to the multimodal nature of agent behaviors and the complex interactions between agents and their environments. Accurately modeling these interactions demands a deep understanding of the environment’s structural constraints, such as road topology, which provide a critical context for forecasting dynamic behaviors.
Standard-definition (SD) maps provide static information for the vast majority of roads worldwide but lack the precision and semantic richness required for advanced tasks such as autonomous driving. Recent studies have explored methods to extract static context from high-definition (HD) maps for trajectory prediction, leveraging autonomous driving datasets that include pre-built HD maps. However, reliance on pre-built HD maps presents critical limitations: their high construction costs and labor-intensive maintenance restrict coverage to limited regions and preclude real-time updates. As a result, trajectory prediction methods that depend on these maps are constrained to mapped areas and struggle to adapt to sudden changes in road conditions, such as construction zones or accidents
For these reasons, methods for constructing local HD maps in real time have been proposed, enabling autonomous vehicles to operate in unmapped areas and rapidly adapt to dynamic road conditions[1, 2, 3]. These real-time HD maps are typically aligned with the ego vehicle’s coordinate frame, helping to mitigate failures caused by localization errors. However, if the prediction module relies solely on the outputs from the map construction process, it becomes highly susceptible to misclassification, missing elements, and positional inaccuracies. Moreover, most map construction methods detect only a limited set of predefined classes (e.g., dividers, boundaries, pedestrian crossings)from raw sensor data, which contains a broad range of environmental information. Without incorporating this additional scene context, an autonomous driving system may overlook critical cues that significantly affect trajectory prediction accuracy.
To better leverage available information, it is important to actively utilize raw sensor data for trajectory prediction, as it contains rich semantic and geometric information that is not fully captured by HD maps. However, sensor data is significantly denser and more redundant than vectorized HD maps, necessitating a mechanism to extract only the most relevant information for trajectory prediction. To address this issue, we propose a novel framework, Bird’s-Eye View Trajectory prediction (BEVTraj). Rather than relying on pre-built HD maps, BEVTraj constructs Bird’s-Eye View (BEV) representations from raw sensor data and directly employs them for trajectory prediction, leveraging deformable attention to extract essential cues from dense BEV features, as illustrated in Fig. 1.
@@
The BEVTraj framework offers several advantages over existing methods. Compared to approaches that rely on HD maps—which provide a high-level abstraction of the driving environment—directly leveraging low-level BEV representations significantly reduces information loss. Additionally, BEVTraj can be seamlessly integrated into modern autonomous driving systems that adopt the BEV paradigm [4, 5, 6]. By incorporating deformable attention, BEVTraj selectively aggregates only the most relevant features for accurate trajectory prediction. Furthermore, we introduce the Sparse Goal Candidate Proposal (SGCP) module to address the issue that the density of goal candidates strongly influences the performance of existing goal-based methods. With this module, BEVTraj supports fully end-to-end trajectory prediction without the need for post-processing steps such as non-maximum suppression (NMS). Our contributions are summarized as follows.
  • We propose Bird’s-Eye View Trajectory Prediction (BEVTraj), a novel goal-driven trajectory prediction framework that directly leverages BEV representations constructed from raw sensor data, thereby eliminating the reliance on pre-built HD maps.
  • BEVTraj utilizes low-level BEV feature representations that effectively capture essential scene context while substantially reducing information loss compared to high-level HD maps.
  • BEVTraj actively leverages deformable attention across multiple stages to selectively aggregate only the information critical for trajectory prediction, enhancing efficiency and prediction accuracy in dense and redundant BEV representations.
  • We introduce the Sparse Goal Candidate Proposal (SGCP) module, which alleviates the sensitivity of goal-based methods to goal candidate density and enables fully end-to-end trajectory prediction without the need for post-processing steps such as non-maximum suppression (NMS).
  • Our framework is fully compatible with modern autonomous driving systems adopting the BEV paradigm, facilitating seamless integration into existing pipelines.
@@ II Related Work
Many studies have proposed trajectory prediction methods that leverage static information from pre-built HD maps. Early methods employed convolutional neural networks (CNNs) to process rasterized map images, capturing spatial features but suffering from information loss and suboptimal spatial encoding. These limitations rendered CNN-based approaches less effective for trajectory prediction compared to vectorized or graph-based representations, which provide more structured encodings of road layouts and traffic constraints [7, 8, 9, 10, 11, 12]. With this shift, Transformer-based models have emerged, utilizing self-attention mechanisms [13] to model complex spatial relationships and improve map feature representations [14, 15, 16, 17, 18, 19]. In particular, HiVT [18] adopts an agent-centric representation, where scene elements are normalized with respect to each agent’s coordinate frame to enhance prediction accuracy. In contrast, QCNet [19] adopts a query-centric design, avoiding per-agent normalization and enabling more efficient multi-agent trajectory prediction.
Some approaches adopt goal-based methods, where multiple goal candidates are generated to predict multimodal future trajectories, as seen in DenseTNT[20] and MTR(Motion Transformer)[15].]. Among them, MTR—most relevant to our work—draws inspiration from Conditional DETR [21] and DAB-DETR [22], assigning initial position embeddings to queries to accelerate training convergence. It also introduces a novel decoder structure that formulates motion prediction as a joint optimization problem of global intention localization and local movement refinement through motion query pairs. Like MTR, R-Pred refines initial trajectory predictions using a Tube-Query Attention-based refinement module [23]. Similarly, BEVTraj adopts an iterative refinement process, enabling more flexible and efficient goal candidate selection.
As mentioned in Section I, pre-built HD maps are costly to construct and often insufficient for real-world autonomous driving, which has motivated research on local HD map construction to automate the mapping process [1, 2, 3]. However, errors in map construction—such as misclassifications or omissions—can degrade the performance of downstream tasks, including trajectory prediction and path planning. To address errors in upstream components such as detection and tracking, recent studies have explored unified frameworks that directly perform decision-making [24, 25, 26, 27]. While some of these models incorporate trajectory prediction, others focus solely on planning, treating prediction as an implicit component rather than a standalone task.
Another line of research adopts a joint perception and prediction approach, aiming to overcome the limitations of traditional map construction and perception pipelines [28, 29, 30] . By integrating multi-sensor data, these methods enhance environmental perception and improve the stability of trajectory prediction, ultimately supporting more reliable planning and decision-making. In a similar spirit, our method directly predicts future trajectories from raw sensor inputs. In contrast to previous approaches that tightly couple instance segmentation with motion prediction, we assume that a perception pipeline for detection and tracking has already been executed to obtain the historical trajectories of surrounding agents. This design approach enables robust, map-free trajectory prediction by decoupling perception and prediction while maintaining compatibility with standard detection and tracking pipelines.
@@ III Method
In this section, we present BEVTraj, a framework that predicts the multimodal future distribution of the target agent’s trajectory. BEVTraj comprises two main modules: Scene Context Encoder, and Iterative Deformable Decoder. The Scene Context Encoder generates scene-level context features by capturing complex interactions within the scene, utilizing the historical states of agents and the BEV features extracted from raw sensor data. Finally, the Iterative Deformable Decoder sequentially predicts multiple goal candidates and their corresponding initial trajectories for the target agent, and refines these predictions iteratively. An overview of BEVTraj is shown in Fig. 2.
@@ III-A Scene Context Encoder
A clear understanding of complex interactions within the scene context is crucial for trajectory prediction. Motion TRansformer (MTR) [15] encodes scene context by utilizing a PointNet-like [31] polyline encoder to vectorize road maps and agent trajectories, applying local self-attention to preserve spatial locality and enhance memory efficiency, and using an auxiliary regression head to predict dense future trajectories of surrounding agents to improve interaction modeling. The Scene Context Encoder builds upon this design and introduces three sub-modules: Sensor Encoder, Pre-Encoder, and BEV Deformable Aggregation (BDA) modules.
@@ III-A1 Sensor Encoder
As in object detection, both semantic and geometric information are essential for trajectory prediction, motivating BEVTraj to adopt a sensor fusion architecture that integrates images and point clouds. This raises a critical question: which spatial domain is best suited for representing essential features? BEVTraj adopts the BEV space for the following reasons. First, since trajectory prediction is conducted in BEV space, this alignment ensures consistency between the prediction output and the feature representation. Moreover, Liu et al. [6] argued that lidar point clouds and camera images are fundamentally different modalities and that fusing them in BEV space is the most appropriate approach. Based on these insights, BEVTraj employs BEVFusion [6]—an off-the-shelf sensor fusion architecture—as the Sensor Encoder to generate a BEV feature map $\mathbf{B}\in\mathbb{R}^{C\times H\times D}$, where H and W denote the spatial dimensions. The BEVFusion module can easily be replaced with any other BEV-based fusion model.
The BEV representation differs from HD maps in several respects. HD maps provide high-level abstractions of static elements (e.g., dividers, boundaries, and pedestrian crossings). In contrast, BEV representations derived from raw sensor data contain redundant and often irrelevant information due to the dense, image-like nature of the BEV grid. Therefore, BEVTraj adopts a strategy that aggregates scene information from $\mathbf{B}$ using deformable attention [32], which dynamically computes offsets to identify key points, enabling the model to selectively extract features essential for trajectory prediction. Sections III-B and III-C provide detailed descriptions of how deformable attention is applied to $\mathbf{B}$ for different purposes.
@@ III-A2 Pre-Encoder
Previous studies categorize scene interaction encoding into two approaches: preserving or compressing the temporal dimension. The former allows for temporal positional encoding and better captures motion tendencies. In contrast, the latter compresses the temporal axis, improving parameter efficiency by increasing representational capacity with the same number of parameters. MTR adopts the latter approach, which limits its ability to model motion dynamics. To address this limitation, the Scene Context Encoder incorporates a Pre-Encoder that applies temporal self-attention (as introduced by Girgis et al. [14]) prior to temporal dimension. enables more effective modeling of motion tendencies before local self-attention is applied. Furthermore, social self-attention (also introduced by Girgis et al. [14]) follows temporal self-attention to capture inter-agent interactions at each timestep. Finally, a PointNet-like encoder transforms the representation from $\mathbb{R}^{N_{a}\times t\times d}$ into an agent feature map $\mathbf{A}\in\mathbb{R}^{N_{a}\times D}$, where $N_{a}$ denotes the number of agents in the scene and $t$ represents the number of past temporal steps.
@@ III-A3 BEV Deformable Aggregation
Unlike MTR, which relies on pre-built HD maps, BEVTraj encodes road topology and physical constraints from raw sensor data into BEV feature map $\mathbf{B}$. In the process of scene context encoding, it is essential to consider not only agent interactions but also these structural constraints. However, due to the high computational cost, global attention—which computes attention scores across all spatial locations—is infeasible on $\mathbf{B}$. To address this, we propose the BEV Deformable Aggregation (BDA) module, which selectively attends to a compact set of key spatial locations in $\mathbf{B}$ and aggregates them into a fixed number of vectors, referred to as the BEV aggregated feature $\in\mathbb{R}^{N_{m}\times D}$. Here, $N_{m}$ denotes the number of selected spatial locations in the BEV feature map used for aggregation, thereby enabling computationally efficient attention operations.
@@
The BEV aggregated feature produced by BDA is then concatenated with the agent feature map $\mathbf{A}$ generated by Pre-Encoder, followed by local self-attention. Subsequently, dense future motion prediction is performed for each agent to account for interactions across future trajectories, resulting in the construction of the scene context feature $\mathbf{C}\in\mathbb{R}^{N_{a}\times D}$.

### v2

@@ Abstract
In autonomous driving, trajectory prediction is essential for safe and efficient navigation. While recent methods often rely on high-definition (HD) maps to provide structured environmental priors, such maps are costly to maintain, geographically limited, and unreliable in dynamic or unmapped scenarios. Directly leveraging raw sensor data in Bird’s-Eye View (BEV) space offers greater flexibility, but BEV features are dense and unstructured, making agent-centric spatial reasoning challenging and computationally inefficient. To address this, we propose Bird’s-Eye View Trajectory Prediction (BEVTraj), a map-free framework that employs deformable attention to adaptively aggregate task-relevant context from sparse locations in dense BEV features. We further introduce a Sparse Goal Candidate Proposal (SGCP) module that predicts a small set of realistic goals, enabling fully end-to-end multimodal forecasting without heuristic post-processing. Extensive experiments show that BEVTraj achieves performance comparable to state-of-the-art HD map-based methods while providing greater robustness and flexibility without relying on pre-built maps. The source code is available at https://github.com/Kongminsang/bevtraj.
@@ I Introduction
Trajectory prediction plays a pivotal role in autonomous driving, enabling vehicles to anticipate the future movements of surrounding agents and make safer, more informed decisions. Beyond individual agent-level forecasting, trajectory prediction also has broader implications for intelligent transportation systems, supporting traffic safety assessment, flow analysis, and risk identification. Recent transportation studies highlight that high-resolution trajectory data and structured scene reconstruction are essential for understanding traffic dynamics and reliably extracting motion patterns from large-scale observations [1, 2, 3], underscoring its role in improving overall traffic efficiency and safety.
Despite these advantages, trajectory prediction remains inherently challenging due to the multimodal nature of agent behaviors and complex agent–environment interactions. Accurate forecasting requires modeling structural constraints such as road topology and scene geometry, for which map information is commonly employed. While standard-definition (SD) maps provide broad but coarse static context, high-definition (HD) maps offer detailed layouts and semantics that better support fine-grained motion reasoning. However, HD maps are costly to construct and maintain, geographically limited, and difficult to update in real time, restricting map-dependent methods to pre-mapped regions and reducing their adaptability to dynamic changes such as construction zones or accidents.
For these reasons, recent works construct local HD maps online from raw sensors, enabling autonomous vehicles to operate in unmapped areas and adapt to dynamic road conditions [4, 5, 6]. However, relying solely on reconstructed maps remains problematic: HD maps are inherently unavailable or unreliable in irregular or rapidly changing environments, and online map construction is still susceptible to perception errors such as misclassification, missing elements, positional inaccuracies, and limited predefined semantics. Consequently, to fully exploit the rich geometric and semantic cues that are not captured by map representations alone and to minimize information loss, trajectory prediction must directly leverage raw sensor data. Fig. 1 contrasts this direct sensor-based paradigm with the conventional map-based approaches.
Despite its importance, directly leveraging raw sensor data introduces new challenges. Unlike vectorized HD maps that provide sparse, semantically structured, and topology-aware representations, sensor-derived Bird’s-Eye View (BEV) features are inherently dense, image-like, and unstructured, requiring the model to jointly reason about both what each feature represents and where it is located, thereby making spatial reasoning substantially more complex. This difficulty is further amplified by the intrinsic nature of trajectory prediction: while perception tasks benefit from globally dense and largely object-agnostic representations, trajectory prediction is fundamentally agent-centric and conditional, where the spatial context relevant to future motion depends on the target agent’s state, behavior mode, and prediction horizon. Consequently, only a small and dynamically varying subset of the scene is informative at each stage, a property that cannot be effectively captured by global attention [7], fixed local-window mechanisms [8], or convolutional aggregation. Naively processing the entire BEV grid is thus both computationally inefficient and prone to diluting critical cues with irrelevant information.
These observations suggest that the central challenge of BEV-based, map-free trajectory prediction is not merely how to encode dense BEV features, but how to selectively and adaptively aggregate task-relevant spatial context in a computationally efficient and agent-specific manner. To address this issue, we propose Bird’s-Eye View Trajectory Prediction (BEVTraj), a framework that performs trajectory prediction through adaptive, agent-centric context aggregation grounded in deformable attention [9]. Specifically, BEVTraj first constructs BEV representations directly from raw sensor data and then selectively gathers context for each agent from a sparse set of dynamically predicted sampling locations instead of exhaustively attending to the entire BEV grid, enabling efficient, geometry-aware, and agent-centric information extraction from dense BEV representations.
@@
In addition, we introduce a Sparse Goal Candidate Proposal (SGCP) module to address a fundamental limitation of existing goal-based approaches, which rely on densely sampled or predefined anchors that are inflexible and often produce redundant or physically implausible candidates. Rather than enumerating dense goals, SGCP directly predicts a small set of sample-adaptive and realistic proposals conditioned on the scene context, enabling efficient and reliable multimodal forecasting with minimal candidates and eliminating the need for heuristic post-processing such as non-maximum suppression (NMS).
@@ II RELATED WORK
II-A Trajectory Prediction with HD Maps
HD-map-based trajectory prediction exploits lane geometry, road topology, and traffic semantics as strong priors to constrain plausible futures. Early methods often rasterize HD maps into BEV images and encode them with CNNs [10, 11, 12, 13, 14]. While rasterization simplifies feature extraction, it can blur lane directionality and break topological continuity, which weakens fine-grained agent–map interaction modeling. To preserve structure, vectorized representations explicitly encode lanes as polylines/graphs with connectivity. Representative works such as VectorNet [15] and LaneGCN [16] model agent–lane and lane–lane relations via hierarchical encoding and graph propagation, and LaneATT [17] further strengthens lane-centric semantic extraction for interaction reasoning.
Beyond map encoding, handling the intrinsic multi-modality of future motion is another key challenge. Goal-conditioned pipelines such as TNT [18] decouple high-level intention (endpoint) from low-level trajectory refinement, and dense goal sampling variants (e.g., DenseTNT) improve coverage of diverse behaviors [18]. More recently, transformer-based architectures become dominant by unifying heterogeneous scene inputs and capturing long-range dependencies, including Wayformer [19], DeMo [20], and query-based decoders such as MTR [21] and SceneTransformer [22]. Extensions like QCNet [23] emphasize query-centric goal reasoning, while FIM [24] incorporates reward-guided inference to inject planning priors. Despite strong accuracy, these paradigms fundamentally rely on the availability and correctness of HD maps; missing lane segments, outdated topology, and domain shifts in map quality can severely degrade prediction and hinder scalability, motivating map-free forecasting.
II-B Map-Free Trajectory Prediction
To reduce the cost and brittleness introduced by HD maps, recent studies explore map-free trajectory prediction [25, 26, 27, 28], where future motion is inferred from observed dynamics and interactions without explicit lane topology. A common line of work treats the problem as generative modeling in trajectory space. For example, DLow [29] learns a latent flow to generate diverse futures and is widely used as a strong map-free baseline. Transformer-based sequence models such as AgentFormer [30] further enhance temporal and inter-agent reasoning within a unified attention framework, improving multi-modal prediction without map priors.
Another effective direction is graph-based interaction modeling, which explicitly represents agents as relational graphs. CRAT-Pred [25] leverages a crystal-graph representation with multi-head attention to strengthen spatial interaction modeling, while GATraj [31] integrates graph neural networks and attention to enrich relational reasoning and achieves competitive performance on nuScenes. Multi-level spatio-temporal designs such as ML-STM [26] reinforce interactions across different temporal scales to better handle complex dynamics. However, many map-free approaches still operate primarily in trajectory/interaction space, where scene geometry and spatial constraints are only implicitly captured. This can be insufficient in topologically complex situations (e.g., merges/splits) or occlusion-heavy scenes, where structured spatial priors are crucial to rule out physically implausible futures.
II-C Map-Free Trajectory Forecasting via BEV Representation
BEV representation provides a geometry-consistent plane for aggregating spatial, temporal, and semantic cues, enabling unified reasoning over scene layout, motion, and interactions. Recent BEV perception frameworks demonstrate scalable feature lifting and projection into BEV [32, 33, 34]. A key driver behind modern BEV transformers is deformable attention [9], which performs sparse sampling to focus computation on geometrically relevant regions while retaining relational modeling capacity. This mechanism underpins architectures such as BEVFormer [35], BEVFormerV2 [36], and Ego3RT [37], enabling effective cross-view fusion and temporal reasoning in dynamic scenes.
Beyond camera-centric pipelines, radar–camera BEV perception has gained attention due to radar’s robustness under adverse weather and long-range sensing capability. Representative frameworks such as RCBEVDet [38] and RCBEVDet++ [39] show that radar reflections can be lifted into BEV and fused with camera features to improve spatial completeness and dynamic cue extraction, suggesting BEV representations naturally extend to multimodal settings.
For forecasting, early BEV predictors constructed rasterized BEV maps from HD-map geometry and processed them with CNNs/MLPs [40, 41, 42], which still inherit HD-map dependency. In contrast, vision-centric frameworks such as FIERY [43], BEVerse [44], and TBP-Former [45] learn BEV dynamics directly from sensors without maps, confirming that BEV can serve as a structured spatial prior for map-free forecasting. Nevertheless, many existing BEV forecasting pipelines remain vision-centric or weakly multimodal, which may limit robustness under adverse sensing conditions. Motivated by this gap, BEVTraj extends deformable-attention BEV modeling toward a fully sensor-based, map-free formulation: it constructs BEV features directly from raw multi-sensor data and jointly models temporal dependencies and inter-agent interactions within a unified geometry-aware representation, enabling robust end-to-end trajectory forecasting under map-free conditions.
@@ III Method
In this section, we present BEVTraj, a framework that predicts the multimodal future distribution of the target agent’s trajectory. BEVTraj comprises two main modules: Scene Context Encoder, and Iterative Deformable Decoder. The Scene Context Encoder generates scene-level context features by capturing complex interactions within the scene, utilizing the historical states of agents and the BEV features extracted from raw sensor data. Subsequently, the Iterative Deformable Decoder sequentially predicts multiple goal candidates and their corresponding initial trajectories for the target agent, and refines these predictions iteratively. An overview of BEVTraj is shown in Fig. 2.
@@ III-A Scene Context Encoder
A clear understanding of complex interactions within the scene context is crucial for trajectory prediction. Motion TRansformer (MTR) [46] encodes scene context by utilizing a PointNet-like [47] polyline encoder to vectorize road maps and agent trajectories, applying local self-attention to preserve spatial locality and enhance memory efficiency, and using an auxiliary regression head to predict dense future trajectories of surrounding agents to improve interaction modeling, which provides an explicit supervision signal for agent-to-agent interaction modeling at the scene (system) level. The Scene Context Encoder builds upon this design and introduces three sub-modules: Sensor Encoder, Pre-Encoder, and BEV Deformable Aggregation (BDA) modules.
@@ III-A1 Sensor Encoder
As in object detection, both semantic and geometric information are essential for trajectory prediction, motivating BEVTraj to adopt a sensor fusion architecture that integrates images and point clouds. This raises a critical question: which spatial domain is best suited for representing essential features? BEVTraj adopts the BEV space for the following reasons. First, since trajectory prediction is conducted in BEV space, this alignment ensures consistency between the prediction output and the feature representation. Moreover, Liu et al. [48] argued that lidar point clouds and camera images are fundamentally different modalities and that fusing them in BEV space is the most appropriate approach. Based on these insights, BEVTraj employs BEVFusion [48]—an off-the-shelf sensor fusion architecture—as the Sensor Encoder to generate a BEV feature map $\mathbf{B}\in\mathbb{R}^{C\times H\times D}$, where H and W denote the spatial dimensions. Since BEVTraj follows the BEV paradigm [49, 50, 48], it can be seamlessly integrated into modern BEV-based autonomous driving systems. Meanwhile, the BEVFusion module serves as an off-the-shelf sensor encoder and can be readily replaced with any alternative BEV-based fusion architecture.
@@ III-A2 Pre-Encoder
Previous studies categorize scene interaction encoding into two approaches: preserving or compressing the temporal dimension. The former allows for temporal positional encoding and better captures motion tendencies. In contrast, the latter compresses the temporal axis, improving parameter efficiency by increasing representational capacity with the same number of parameters. MTR adopts the latter approach, which limits its ability to model motion dynamics. To address this limitation, the Scene Context Encoder incorporates a Pre-Encoder that applies temporal self-attention (as introduced by Girgis et al. [51]) prior to temporal dimension. enables more effective modeling of motion tendencies before local self-attention is applied. Furthermore, social self-attention (also introduced by Girgis et al. [51]) follows temporal self-attention to capture inter-agent interactions at each timestep. Finally, a PointNet-like encoder transforms the representation from $\mathbb{R}^{N_{a}\times t\times d}$ into an agent feature $\mathbf{A}\in\mathbb{R}^{N_{a}\times D}$, where $N_{a}$ denotes the number of agents in the scene and $t$ represents the number of past temporal steps.
@@ III-A3 BEV Deformable Aggregation
As described in Section I, sensor-derived Bird’s-Eye View (BEV) features are inherently dense, image-like, and unstructured. BEVTraj encodes road topology and physical constraints directly from raw sensor data into a BEV feature map $\mathbf{B}$, which makes global attention—computing attention scores across the entire BEV grid—computationally prohibitive. To address this challenge, we propose the BEV Deformable Aggregation (BDA) module, which selectively attends to a compact set of key spatial locations in $\mathbf{B}$ and aggregates them into a fixed number of vectors, referred to as the BEV aggregated feature $\in\mathbb{R}^{N_{m}\times D}$. Here, $N_{m}$ denotes the number of selected spatial locations used for aggregation, enabling computationally efficient attention operations.
@@
The BEV aggregated feature produced by BDA is then concatenated with the agent feature $\mathbf{A}$ generated by Pre-Encoder, followed by local self-attention to model interactions between surrounding agents and the environment. Subsequently, dense future motion prediction is performed for each agent to account for agent interactions over the future horizon, resulting in the construction of the scene context feature $\mathbf{C}\in\mathbb{R}^{N_{a}\times D}$.